OpenStudy (ttop0816):

is my answer correct?? Given that θ terminates in Quadrant II and sinθ = 12/13 , find secθ. -13/5 <--- i chose this! -13/12 13/12 13/5

OpenStudy (anonymous):

yeah, y r right! cuz cosine(theta)=-5/13 then sec=1/cos=-13/5

OpenStudy (kainui):

I thought this was the same question! Haha yeah right!

OpenStudy (anonymous):

u can draw a unit circle to make it clearer if u encounter the problem like that next time! It is a quite good method.